What are the first three terms of the arithmetic sequence if the 25th term is 35 and the 30th term is 5.?
Answer by stanbon(75887)

What are the first three terms of the arithmetic sequence if the 25th term is 35 and the 30th term is 5.?
-------------------------
a(25) = a(1) + 24d = 35
a(30) = a(1) + 29d = 5
---------------------------
Subtract and solve for "d":
5d = -30
d = -6
------------------
Solve for a(1)::
a(1) + 24(-6) = 35
a(1) + -144 = 35
a(1) = 179
------
a(2) = 179 -6 = 173
a(3) = 179 + 2(-6) = 167
---------------
Cheers,
Stan H.